OpenStudy (anonymous):

Simplify 1/2x^2 + 3/3y - 1/2y

OpenStudy (anonymous):

\[\frac{ 1 }{ 2x ^{2} } + \frac{ 3 }{ 3y } - \frac{ 1 }{ 2y }\]

OpenStudy (anonymous):

(x^2+y)/(2 x^2 y)